OpenStudy (annie12m):

Is the set of numbers that can be written as the product of 6 and an integer closed under subtraction? A. Yes, because the difference of any two multiples of 6 is equal to 6. B. Yes, because the difference of any two multiples of 6 is also a multiple of 6. C. No, and a counterexample is 6 – 6 = 0. D. No, and a counterexample is 6 – 18 = –12.

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @annie12m @~ You might not have noticed but you are in the "OpenStudy Feedback" subject. Hover your mouse where it says OpenStudy Feedback and find the correct subject by clicking the 'Find More Subjects' button or simply clicking the correct subject in the list that you have already studied! Alternately, you can get to the correct subject by going to this link: http://openstudy.com/study#/groups/Mathematics

OpenStudy (samsungfanboy):

Welcome to OpenStudy @annie12m . As you see, the Greatest OS user of all time has just commented on you're post. @TheSmartOne Mostly known as, The White Mamba. Be sure to fan him and give him a medal and as always check his amazing legendary tutorial just in case you get lost. If you have an questions, you can always ask the Professor here. Thankyou! And again, welcome to OpenStudy (: ~SamsungFanBoy aka The Mixed Mamba

OpenStudy (shawnscientifica):

So, here is my logic. I begin by analyzing the question, they want to know if 6x - 6y is within the domain of 6D. Meaning is the difference between any two multiples of 6 able to be rewritten as 6D where you can make D whatever you want. Even Pizza (just kidding, but you get it.) and since 6x-6y = 6(x-y) which (x-y) can be made into another variable like... Maybe.... D? Then you CAN rewrite 6x-6y = 6(x-y) = 6D where D = (x-y) then it's true! You can write any difference between two multiples of 6 as 6D and so it's closed under Subtraction. When something is closed under something it means when you do that operation (subtraction in this case) it doesn't change the original formula (which in this case was 6x and it doesn't since 6x can equal 6D since we can make D whatever we want).